# Mitigating the Alignment Tax of RLHF

### v3

@@ Appendix C Implementation Details
C.1 Rejection Sampling Fine-tuning Implementation
@@ C.2 Implementation of PPO
First, we follow (Ramamurthy et al., 2022) to warm up by finetuning the model on the preferred samples of the preference dataset for 1 epoch for a more stable training process. Moreover, in contrast to the implementation in traditional DRL scenario, for alignment of LLMs, following (Ziegler et al., 2019; Wu et al., 2021; Ouyang et al., 2022; Rafailov et al., 2023; Liu et al., 2023), we will also modify the reward optimization as the following KL-regularized form:
@@ C.3 Implementation of DPO
C.4 Implementations of Existing Methods to Alleviate Alignment Tax
We test existing methods mainly on the RSF method which is implemented as discussed in Appendix C.1. Details about how we implement existing methods to mitigate forgetting are described as follows.
@@ C.4 Implementations of Existing Methods to Alleviate Alignment Tax
C.5 Implementations of Adaptive Model Averaging
@@
Appendix D More Results
D.1 The Alignment Tax during Training (Results of Early Stopping)
@@
Figure 12: The reward-tax trade-off during training
D.2 More Results of Averaging Different Parts
@@
D.3 Comparison of RLHF Algorithms
We compare the reward-tax trade-off of RSF, DPO and PPO in Figure 14. We observe that RSF is consistently better than DPO. However, we also note that this is not a fair comparison since DPO does not directly optimize for the reward.
Figure 14: Comparison of RLHF algorithms in terms of reward-tax trade-off.
D.4 Detailed Results of Adaptive Model Averaging
We provide the detailed results of adaptive model averaging on various benchmarks, e.g., Reading Comprehension, Commonsense QA and translation, and different RLHF methods, e.g., RSF, PPO, and DPO.
Figure 15: Detailed results of adaptive model averaging on various benchmarks and RLHF methods.
Appendix E Theoretical Settings, Proofs and Discussions
E.1 Re-statement of Formal Settings
@@
Definition E.1 (Data Generation Process).
@@
Definition E.2 (Model Averaging, Definition 4 of (Lin et al., 2023)).
@@
Assumption E.3 (Small Noise).
@@
Assumption E.4 (Orthogonal features (Lin et al., 2023; Allen-Zhu and Li, 2020)).
@@
We first estimate the accuracy of $f_{\mathrm{avg}}$ on task ($a$), i.e., $\mathcal{A}_{a}(f_{\mathrm{avg}})$, for a sample from class $k\in 1,\cdots,K$ and $k^{\prime}\neq k,k^{\prime}\in 1,\cdots,K$. Then by $|\mathcal{Y}_{a}\cap\mathcal{Y}_{b}|=0$ and Assumption E.4, we have
@@
E.3 Discussion on the Effect of Task Similarity on Model Averaging
@@
Appendix F Hyper-Parameters
@@
Table 3: Hyper-parameters for AMA experiments.
@@
DPO AMA Scheduler Cosine decay with 0.03 warm-up

### v4

@@ Appendix D Implementation Details
D.1 Rejection Sampling Fine-tuning Implementation
@@ D.2 Implementation of PPO
First, we follow (Ramamurthy et al., 2022) to warm up by finetuning the model on the preferred samples of the preference dataset for 1 epoch for a more stable training process. Moreover, in contrast to the implementation in traditional DRL scenario, for alignment of LLMs, following (Ziegler et al., 2019; Wu et al., 2021a; Ouyang et al., 2022; Rafailov et al., 2023; Liu et al., 2023), we will also modify the reward optimization as the following KL-regularized form:
@@ D.3 Implementation of DPO
D.4 Implementations of Existing Methods to Alleviate Alignment Tax
We test existing methods mainly on the RSF method which is implemented as discussed in Appendix D.1. Details about how we implement existing methods to mitigate forgetting are described as follows.
@@ D.4 Implementations of Existing Methods to Alleviate Alignment Tax
D.5 Implementations of Heterogeneous Model Averaging
@@
Appendix E More Results
E.1 The Alignment Tax during Training (Results of Early Stopping)
@@
Figure 12: The alignment-forgetting trade-off during training
E.2 More Results of Averaging Different Parts
@@
E.3 Comparison of RLHF Algorithms
We compare the alignment-forgetting trade-off of RSF, DPO and PPO in Figure 14. We observe that RSF is consistently better than DPO. However, we also note that this is not a fair comparison since DPO does not directly optimize for the reward.
Figure 14: Comparison of RLHF algorithms in terms of alignment-forgetting trade-off.
E.4 Results of AdaMerging Yang et al. (2023)
Previous studies Yang et al. (2023) have also discussed the idea of dynamically assigning different weights to different layers when merging models, aiming to maximize performance on a specific task (e.g., $\mathcal{T}_{i}$). These approaches assume access to the task-specific data $\mathcal{T}_{i}$. However, considering the nature of alleviating alignment tax, which aims to mitigate forgetting across a extremely wide range of tasks ($\mathcal{T}_{j_{1}}...\mathcal{T}_{j_{K}}$), these methods fail to effectively optimize performance for multiple tasks simultaneously. Specifically, we want to preserve the abilities on a wide range of tasks and it is hard to get the data for all these tasks. Further more, some ability such as in-context learning does not have a clear corresponding training set. So it is less practical to find training set for AdaMerging.
Here we demonstrate when we use AdaMerging to optimizes for task A and the training set does not cover task B, AdaMerging can not preserve the ability on task B. Specifically, we provide AdaMerging with labeled data for Reading Comprehension (i.e., task A) and optimize the 26 layer-wise merging ratios as Yang et al. (2023). To have a clear comparison with vanilla model averaging, we try different mean averaging ratio for AdaMerging among 0.2, 0.4 and 0.6. We also show both the results on task A and B.
Figure 15: Results of AdaMerging. We optimize AdaMerging on Reading Comprehension and found it can hardly do well on Common Sense.
In contrast, our HMA only require the RLHF data and does not need any data from the tasks which we want to preserve ability. Figure 16 shows that HMA can alleviate the alignment tax evaluated on a wide range of tasks.
E.5 Detailed Results of Heterogeneous Model Averaging
We provide the detailed results of Heterogeneous model averaging on various benchmarks, e.g., Reading Comprehension, Commonsense QA and translation, and different RLHF methods, e.g., RSF, PPO, and DPO.
Figure 16: Detailed results of Heterogeneous model averaging on various benchmarks and RLHF methods.
Appendix F Theoretical Settings, Proofs and Discussions
F.1 Re-statement of Formal Settings
@@
Definition F.1 (Data Generation Process).
@@
Definition F.2 (Model Averaging, Definition 4 of (Lin et al., 2023)).
@@
Assumption F.3 (Small Noise).
@@
Assumption F.4 (Orthogonal features (Lin et al., 2023; Allen-Zhu and Li, 2020)).
@@
We first estimate the accuracy of $f_{\mathrm{avg}}$ on task ($a$), i.e., $\mathcal{A}_{a}(f_{\mathrm{avg}})$, for a sample from class $k\in 1,\cdots,K$ and $k^{\prime}\neq k,k^{\prime}\in 1,\cdots,K$. Then by $|\mathcal{Y}_{a}\cap\mathcal{Y}_{b}|=0$ and Assumption F.4, we have
@@
F.3 Discussion on the Effect of Task Similarity on Model Averaging
@@
F.4 Close Form of $F_{p}(x)$
Here we provide the explicit expression of function $F_{p}(x)$ in $K$ class situation, which is monotonically increasing with $x$.
We denote a $K-1$-dim random variable $\boldsymbol{\eta}\sim\mathcal{N}(\boldsymbol{x},\boldsymbol{M})$, in which
$\boldsymbol{M}_{i,i}=\frac{p(K+2-pK)}{K},\boldsymbol{M}_{i,j}=\frac{p(K+1-pK)}{K},$
then $F_{p}(x)$ is defined as
$F_{p}(x)=\mathbb{P}(\boldsymbol{\eta}_{1}>0,\dots,\boldsymbol{\eta}_{K-1}>0).$
Appendix G Hyper-Parameters
@@
Table 4: Hyper-parameters for HMA experiments.
@@
DPO HMA Scheduler Cosine decay with 0.03 warm-up